### v1

# CATD: Unified Representation Learning for EEG-to-fMRI Cross-Modal Generation

Abstract
Multi-modal neuroimaging analysis is crucial for a comprehensive understanding of brain function and pathology, as it allows for the integration of different imaging techniques, thus overcoming the limitations of individual modalities. However, the high costs and limited availability of certain modalities pose significant challenges. To address these issues, this paper proposed the Condition-Aligned Temporal Diffusion (CATD) framework for end-to-end cross-modal synthesis of neuroimaging, enabling the generation of functional magnetic resonance imaging (fMRI)-detected Blood Oxygen Level Dependent (BOLD) signals from more accessible Electroencephalography (EEG) signals. By constructing Conditionally Aligned Block (CAB), heterogeneous neuroimages are aligned into a potential space, achieving a unified representation that provides the foundation for cross-modal transformation in neuroimaging. The combination with the constructed Dynamic Time-Frequency Segmentation (DTFS) module also enables the use of EEG signals to improve the temporal resolution of BOLD signals, thus augmenting the capture of the dynamic details of the brain. Experimental validation demonstrated the effectiveness of the framework in improving the accuracy of neural activity prediction, identifying abnormal brain regions, and enhancing the temporal resolution of BOLD signals. The proposed framework establishes a new paradigm for cross-modal synthesis of neuroimaging by unifying heterogeneous neuroimaging data into a potential representation space, showing promise in medical applications such as improving Parkinson’s disease prediction and identifying abnormal brain regions.
{IEEEkeywords}
Cross-Modal Generation, Representation Learning, Diffusion Model, Functional Neuroimaging, Temporal Super-Resolution
1 Introduction
[…]
The BOLD signal, measured by fMRI, provides a detailed and precise mapping of brain activity[1]. The sensitivity of this signal to changes in oxygenation and deoxygenation levels in the blood provides a dynamic image of brain function. This helps to understand and track various brain diseases and is even considered the gold standard of modern functional neuroimaging[2]. Notably, BOLD fMRI has provided important insights into the brain dynamics of disorders such as ischemic stroke[3], schizophrenia[4], Alzheimer’s disease[5], focal epilepsy[6], and depression[7]. This highlights its indispensable role in diagnosing and monitoring these diseases.
Although BOLD fMRI is highly regarded for its detailed imaging capabilities, acquiring such scans is not only costly[8] and time-consuming[9], but also has limitations in a variety of clinical setting applications[10, 11]. At the same time, the inherent delay between neuronal activity and haemodynamic responses captured by BOLD fMRI can limit its effectiveness in real-time brain activity monitoring[12]. In contrast, EEG has several advantages in reflecting brain activity. Because EEG directly measures brain electrical activity, it provides real-time insight with high temporal resolution, which is critical for capturing rapid dynamic changes in neuronal circuits[13]. However, the lower spatial resolution of EEG limits its ability to accurately localise activity in brain regions[14], and Fig.1 outlines the respective strengths and limitations of fMRI and EEG.
Figure 1: Comparison of the respective advantages and disadvantages of BOLD fMRI and EEG
In recent years, cross-modal neuroimage synthesis has gradually become a research hotspot with the rapid development of artificial intelligence technology[15], especially with the wide application of generative models in different fields. Generative AI models, such as transformer[16], diffusion models[17, 18, 19] and generative adversarial networks (GANs)[20, 21, 22], demonstrate great potential in cross-modal data synthesis[23, 24]. For example, Yan et al. used a deep convolutional generative adversarial network (DCGAN) to reconstruct missing BOLD signals for individual participants[25]. The success of generative AI is based on the existence of some of the same underlying information and potentially relevant features in different modal information. Although EEG and fMRI are acquired in different ways, they both reflect brain activity and EEG has the advantage of low cost and fewer limitations on its use. It has been found that there is a strong correlation between microstate transitions in EEG signals and BOLD signals[26] and that BOLD functional connectivity correlates with functional connectivity of EEG activity[27]. By simultaneously recording EEG and fMRI signals, researchers can observe the relationship between EEG activity and metabolic activity in the brain. When electrical brain activity increases, blood oxygen levels in the brain also increase, suggesting a one-to-one correspondence between changes in electrical brain activity and the state of metabolic activity in the brain[28, 29]. This correlation provides theoretical possibilities for cross-modal reconstruction and temporal resolution enhancement of fMRI using generative AI and EEG.
EEG and BOLD signals are highly heterogeneous time-series data, requiring powerful generative AI models for cross-modal synthesis. Currently, diffusion transformer model[30] has demonstrated excellent performance in several areas such as image generation. The emergence of Sora[31] further demonstrates the potential application of diffusion models in this direction of temporal data processing. Drawing inspiration from these advancements, the CATD framework for the unified representation of EEG and BOLD signals is proposed. This framework is the first to achieve cross-modal synthesis of high-dimensional, heterogeneous brain functional data using a diffusion model. The novelties and contributions of this paper can be summarized in the following points:
  • (i)
    A new paradigm based on generative AI for unified representation of neuroimaging is proposed. As far as we know, it is the first time a diffusion-driven end-to-end framework is developed for EEG-to-fMRI synthesis. Utilizing low-cost, accessible EEG signals, the proposed CATD framework is capable to synthesis high-cost, difficult-to-access BOLD signals. The proposed framework enables high-quality and stable cross-modal generation from EEG to BOLD signals, bridging the gap between different neuroimaging modalities.
  • (ii)
    The CAB module is designed to align high-temporal, low-spatial resolution EEG signals with low-temporal, high-spatial resolution BOLD signals within a latent space, facilitating a unified representation across modalities. The combination with DTFS module also leverages EEG’s superior temporal resolution to improve the temporal super-resolution of BOLD signals, capturing detailed brain dynamics that surpass traditional methods.
2 Method
2.1 Overview
BOLD signals are valuable in reflecting brain activity and are essential for analysing, diagnosing and treating brain disorders. However, BOLD signal acquisition is not possible in some patients due to medical conditions or other limitations. In response, the paper proposed the CATD framework, an innovative EEG-to-BOLD signal conversion model based on Scalable Diffusion Models with Transformers (DiT)[30]. This approach addresses the limitations of cross-modal reconstruction and temporal super-resolution of whole-brain BOLD signals, which are not possible with existing techniques. The CATD framework addresses the challenges posed by high dimensionality and asymmetry between EEG and BOLD data through a novel heterogeneous alignment method that facilitates dimensional matching and enhances signal compatibility. At the same time, it employs a cross-attention mechanism to efficiently generate cross-modal EEG-modulated BOLD signals. The proposed DTFS achieves the control of the EEG sampling rate by sliding the sampler, which in turn achieves the enhancement of the temporal resolution of the output BOLD signal. Fig.2 depicts the complete structure and functionality of the CATD framework.
Figure 2: The overall framework of the proposed CATD. The upper part of the figure shows how two different dimensions of data are processed to achieve the initial alignment. The lower half shows the reconstruction pipeline of the BOLD signal under the control of the EEG condition based on the DiT structure.
2.2 Data Alignment Method
EEG and BOLD signals are high-dimensional signals, one with high temporal resolution and the other with high spatial resolution, highly heterogeneous in scale, and both need to be processed to the same scale using the following methods to perform integration operations. For preprocessing BOLD signals, we followed the method by Yan et al. [25], utilizing freesurfer[32] software to register brain BOLD activity to the cerebral cortex. The data is then aligned using the fsaverage template and downsampled to the fsaverage4 template, creating a functional map of cortical BOLD activity. These signals are fed into the encoder of a pretrained Variational Autoencoder (VAE) [33] for dimensionality reduction. The reduced data is segmented into patches and transformed into input tokens via an embedding layer, producing the initial state $x_{0}$ for the diffusion model training.
EEG signals, corresponding to the first 6 seconds of each BOLD functional map, are selected due to the 6-second delay between neuronal activity and blood oxygenation response [34, 35]. Feature extraction and dimensionality reduction are performed using a dynamic temporal-frequency analysis method based on the short-time Fourier transform. The reduced EEG data is then divided into segments and converted into tokens that match the BOLD frames in dimension and number through the EEG embedding layer. These tokens serve as the conditional signals $c$ for the diffusion model, enabling alignment of the high-dimensional, heterogeneous time-series EEG data with BOLD signals.
For the BOLD signal super-resolution task, the sample rate of the EEG signal is controlled using the DTFS module. During training, overlapping EEG signal samples are employed. Specifically, the EEG signal is sampled at one-third of its original interval while maintaining the same overall sampling duration. This fine-grained temporal segmentation approach allows the reconstructed BOLD signal to achieve three times the temporal resolution of the original signal. As a result, this method generates more refined BOLD signal sequences, significantly enhancing temporal resolution and providing a more detailed representation of brain activity.
2.3 EEG to BOLD Diffusion
{mdframed}[backgroundcolor=bottomcolor,rightline=false,leftline=false,topline=false,bottomline=false,innerleftmargin=5pt,innerrightmargin=5pt,userdefinedwidth=0.92nerbottommargin=5pt,innertopmargin=5pt]
Define:
$F_{BOLD}(x)$: Function to preprocess BOLD data
$F_{EEG}(c,f_{s})$: DTFS process, with $f_{s}$ as sample rate
Encoder$(x)$: Encoder for fMRI data
Decoder$(x)$: Decoder for fMRI data
CAB$(c)$: Condition-Aligned Block function
$\alpha_{t},\beta_{t}$: Noise schedule parameters
$\epsilon_{\theta}$: Neural network for predicting noise
$\mathcal{L}(\theta)$: Loss function
Input:
$x$ ;
  // Raw fMRI data
$c$ ;
  // Raw EEG data
$f_{s}$ ;
  // EEG sample rate
$N$ ;
  // Number of diffusion steps
Get BOLD Data:
$x\leftarrow$ Load raw fMRI data
$x\leftarrow F_{BOLD}\left(x\right)$
$x\leftarrow Encoder\left(x\right)$
Initialize Diffusion Model:
$x_{0}\leftarrow x$ ;
  // Initialize with preprocessed BOLD data
Forward Diffusion Process:
for $t\leftarrow 1$ to $N$ do
       Sample $\epsilon_{t}\sim\mathcal{N}(0,I)$ ;
       $x_{t}\leftarrow\sqrt{\alpha_{t}}x_{t-1}+\sqrt{\beta_{t}}\epsilon_{t}$
end for
Reverse Denoising Process:
for $t\leftarrow N$ to $1$ do
       $\hat{\epsilon_{t}}\leftarrow\epsilon_{\theta}(x_{t},CAB(c))$ ;
       $x_{t-1}\leftarrow\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon_{t}}\right)$
end for
Loss Calculation and Backpropagation:
$\mathcal{L}(\theta)={\left\|\epsilon_{t}-\epsilon_{\theta}(x_{t},CAB(c))\right\|}^{2}_{2}$
Update model parameters using backpropagation
Algorithm 1 EEG to BOLD Signal Generation using Diffusion Model: Training Phase
{mdframed}[backgroundcolor=bottomcolor,rightline=false,leftline=false,topline=false,bottomline=false,innerleftmargin=5pt,innerrightmargin=5pt,userdefinedwidth=0.92nerbottommargin=5pt,innertopmargin=5pt]
Initialize Diffusion Model for Inference:
$x_{T}\leftarrow Initializewithnoise$
$N\leftarrow Setnumberofdiffusionsteps$
Reverse Denoising Process:
for $t\leftarrow N$ to $1$ do
       $\hat{\epsilon_{t}}\leftarrow\epsilon_{\theta}(x_{t},CAB(c))$ ;
       $x_{t-1}\leftarrow\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon_{t}}\right)$
end for
Decode Reconstructed Signal:
$x\leftarrow Decoder(x_{0})$ ;
  // Decode reconstructed BOLD signal
Output:
Reconstruct final BOLD signal from $x$ ;
Algorithm 2 EEG to BOLD Signal Generation using Diffusion Model: Inference Phase
2.3.1 Basic ideas
As one of the highest performing generative AI models available, diffusion models are known for their stable training process and superior quality of generated output. These models work by gradually transforming the data distribution into a Gaussian distribution through a forward process, and then learning the reverse transformation to generate new data samples. In the forward phase, the diffusion model starts with the original data $x_{0}$ and gradually increases the noise over a number of time steps, eventually reaching an almost completely random state $x_{T}$, a process that can be described by the following Gaussian process:
$q\left(x_{t}|x_{0}\right)=\mathcal{N}\left(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},\left(1-\bar{\alpha}_{t}\right)I\right),$ (1)
where $\bar{\alpha}_{t}$ is a hyperparameter that decreases with time. This process is characterised by a Markov chain and each step is usually controlled by a variance preserving transformation.
The inverse process aims to reconstruct the original data from the noisy state. This is done by training a neural network to predict the noise added at each step of the forward process, and then iteratively removing this noise to recover the original data $x_{0}$. This inverse process can be described by the following equation:
$p_{\theta}\left(x_{t-1}|x_{t},c\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},c\right),\Sigma_{\theta}\left(x_{t},c\right)\right),$ (2)
where $\mu_{\theta}\left(x_{t},c\right)$ and $\Sigma_{\theta}\left(x_{t},c\right)$ are predicted by a neural network with parameter $\theta$. We have adopted a transformer architecture for this network because of its flexibility and excellent ability to capture long-range dependencies, which makes it particularly suitable for processing time-series signals such as EEG. The inherent attentional mechanism of the transformer allows for the integration of conditional information denoted by $c$ through a cross-attentional mechanism. This integration guides the generation process, thus enhancing the applicability of the model to cross-modal generation tasks involving high-dimensional, heterogeneous data.
2.3.2 Architectures
As shown in Fig.2, our proposed EEG to BOLD diffusion model consists of a prediction network that consists of CAB, several layers of Transformer Blocks connected in series, an input embedding layer, and an output section. Each Transformer Block contains mainly one layer of multi-head self-attention mechanism and one layer of multi-head cross-attention mechanism. Among them, the cross-attention layer is responsible for incorporating the condition information into the network. The CAB integrates the EEG markers as condition $c$ as well as the diffusion step $t,t\in\left\{0,1,\dots,T\right\}$, and delivers this integrated conditional information to the cross-attention network of each Transformer Module as a way to achieve an effective input of conditional information.
2.3.3 Loss Function
According to Eq.1 and Eq.2, the inverse process is trained using a log-likelihood variational lower bound on $x_{0}$, which can be simplified as
$\begin{split}\mathcal{L}\left(\theta\right)=&-p\left(x_{0}|x_{1}\right)+\\ &{\textstyle\sum_{t}\mathcal{D}_{KL}\left(q^{*}\left(x_{t-1}|x_{t},x_{0}\right)\parallel p_{\theta}\left(x_{t-1}|x_{t},c\right)\right)},\end{split}$ (3)
where $q^{*}$ denotes is the true conditional distribution of $x_{t-1}$ given $x_{t}$ and $x_{0}$, and $\mathcal{D}_{KL}$ denotes the KL dispersion of the two distributions.
In discussing the training process of the diffusion model, it is crucial to consider the consistency of the entire reconstruction process, which requires the model to complete the computation of all time steps $T$ before each parameter update. While this approach ensures the comprehensiveness of the learning process, it has a significant negative impact on the convergence speed, stability and optimisation efficiency of the model. At the same time, it also places a high demand on computational resources. In view of this, we decided to simplify Eq.3. Instead of seeking to minimise the KL scatter in terms of distributional similarity across all time steps, we instead turned to minimising the prediction error in each time step. This approach not only simplifies the computational process but also helps to improve the training efficiency and stability of the model. The expression of the simplified loss function is as follows:
$\mathcal{L}\left(\theta\right)={\left\|\epsilon_{t}-\epsilon_{\theta}\left(x_{t},CAB(c)\right)\right\|}^{2}_{2},$ (4)
where $\epsilon_{\theta}\left(x_{t},CAB(c)\right)$ denotes the noise predicted by the network and $\epsilon_{t}$ denotes the ground truth sampled Gaussian noise.
3 Experiment
3.1 Experiment Settings
Figure 3: (a) Radar plot of the real BOLD signal, synthetic BOLD signal, real + synthetic BOLD signal, and the results of the ablation of the conditioned blocks in the motion imagery state and the resting state classification results. (b) Radar plot of different frequency bands of EEG synthesized BOLD signals in the motion and resting states after our proposed CATD framework. (c) Radar plot of real EEG signals synthesized BOLD signals, and the combination of the two for the prediction of Parkinson’s disease in a medical decision support experiment.
3.1.1 Dataset
The experiments were performed on the following datasets: the motor imagery dataset[36], the NODDI dataset[37, 38], and the EEG dataset of Parkinsonian patients[39]. For evaluating the reconstruction results, the XP1 part of the motor imagery dataset was used, which contained data from 10 subjects (2 females and 8 males, mean age: 28.4 years $\pm$ 10.6 years). Their paired 64-channel EEG and whole-brain fMRI scans were acquired simultaneously using a block design, with each block consisting of a 20-second rest and a 20-second motor imagery. The motor imagery task contained 48 blocks per subject during the test duration. The NODDI dataset included simultaneous resting-state 64-channel EEG and whole-brain fMRI scan data from 17 adult volunteers (11 males and 6 females, mean age: 32.84 $\pm$ 8.13 years). The EEG dataset of the Parkinson’s patients contained 64-channel resting-state brain EEG data from 8 subjects (4 males and 4 females, mean age: 74.25 years $\pm$ 8.75 years). EEG data in all datasets were obtained using the international 10-20 lead system.
3.1.2 Implementation Detail
The experiments were conducted on a server platform equipped with two Nvidia Tesla A800 compute cards. For model parameters, the depth of the Transformer Block was set to 12, the hidden space dimension of the patch was 768, and the number of heads in the attention network was 12. Training was performed using the Adam optimizer with an initial learning rate of 0.0001, a batch size of 8, and over 1000 epochs. To compute the experimental classification metrics, a five-fold cross-validation method was employed to ensure the reliability of the results. For calculating other quantitative metrics, five independent experiments were conducted, and the results were averaged to ensure data accuracy and stability.
3.1.3 Metrics
A variety of categorical metrics were employed, including Accuracy (ACC), Precision (PRE), Sensitivity (SEN), and F1-score, to demonstrate the performance of the synthesized results in downstream tasks. Additionally, to evaluate reconstruction quality in the spatial dimension, Root Mean Square Error (RMSE) and Structural Similarity Index (SSIM) were used. For the temporal dimension, Cosine Similarity and Concordance Correlation Coefficient (CCC) were used to assess reconstruction quality. For the temporal super-resolution experiments, Signal-to-Noise Ratio (SNR) was also used to evaluate the effectiveness of the synthesized signal, with SNR values being base-10 logarithms. To present the results more intuitively, graphical representations of classification and reconstruction performance in different experiments were provided. For example, in the medical decision support experiment, the potential application of the method in medical tasks was illustrated through difference maps of BOLD function maps.
3.2 Evaluation of the reconstructed BOLD signal
To assess the effectiveness of the EEG-to-BOLD signal reconstruction, both synthetic and real BOLD signals were utilized to differentiate between subjects’ performance in motor imagery and resting states. The brain activity of the subjects was divided into 20-second chunks for both resting and motor imagery tasks. A clear distinction between these states by the model indicated effective learning of brain activity patterns. As shown in Fig.3(a), the cross-modal synthetic BOLD signals outperformed the real data in accuracy and precision. When combined with real data, the classification metrics improved significantly, although recall remained similar to the real data. Ablation experiments demonstrated that CAB enhances the model’s ability to use EEG features to constrain BOLD signal synthesis, improving classification metrics compared to models without the CAB.
Spatial (RMSE, SSIM) and temporal (cosine similarity, CCC) metrics of the synthesized BOLD signal were also calculated for the three motion picture states. The results in Fig.4 show low RMSEs (all below 0.1) and high structural similarity indices (all above 0.6735). The cosine similarity is approximately 0.85, and the CCC value is about 0.8, indicating a high spatio-temporal correlation with the real signal. The ablation experiments (e.g., Fig.3(a) and the No CAB section in Fig.4) confirm the effectiveness of the CAB in improving signal quality.
Figure 4: Results of quantitative spatial and temporal metrics for the synthetic BOLD signal in three states of motor imagery. The upper half shows the metrics obtained using the CATD framework we propose, and the lower half shows the metrics obtained after ablation experiments on conditioned blocks.
To further illustrate the capability of the proposed model in learning high-dimensional heterogeneous brain activity features and achieving cross-modal transitions, t-SNE plots were used (Fig.5). The distribution of BOLD signals synthesized with the full CATD framework more closely matches the real signal distribution, underscoring the critical role of the CAB in enhancing signal quality.
Figure 5: (a) Shows the tSNE plots of the distribution of the synthesized BOLD signal versus the real signal in the CATD framework for the three motor imagery states. (b) Demonstrates the t-SNE plot of the synthetic BOLD signal versus the real signal distribution after ablation of the conditioned block.
3.3 Evaluation of temporal resolution enhanced BOLD signals
To verify that the CATD framework can leverage the high temporal resolution of EEG signals to achieve the temporal super-resolution of BOLD signals, temporal resolution enhancement experiments were conducted. The proposed DTFS was used for EEG to achieve triple temporal super-resolution, meaning that the temporal resolution of the reconstructed BOLD signal was three times that of the actual BOLD signal. Since the application scenario for temporal super-resolution typically involves enhancing the existing BOLD signal rather than generating it in the absence of a BOLD signal, the original low temporal resolution BOLD signal was used as a known condition. The enhanced high temporal resolution BOLD signal was obtained by constraining the generated signal using the low-resolution original BOLD signal. The hyper-temporal resolution reconstruction results were similarly evaluated in three different motor imagery states, as shown in Table 1.
Table 1: Results of the time series similarity between the reconstructed triple time-resolved BOLD signal and the real BOLD signal in three motion imagery states.
Cosine Similarity CCC
eegNF 0.98998$\pm$0.00008 0.98589$\pm$0.00003
fmriNF 0.99187$\pm$0.00007 0.98714$\pm$0.00006
eegfmriNF 0.98740$\pm$0.00005 0.98209$\pm$0.00002
Table 2: Comparison of SNR between real BOLD signals with low temporal resolution and synthesized BOLD signals with high temporal resolution in three motion imagery states.
Real Synthetic
eegNF 12.9640 12.9995$\pm$0.0103
fmriNF 13.4383 13.3112$\pm$0.0137
eegfmriNF 12.4715 13.0708$\pm$0.0533
An important advantage of high temporal resolution is the potential for signal-to-noise ratio (SNR) enhancement. A higher SNR implies a higher proportion of useful information in the signal relative to the noise and signals with a higher SNR can more reliably reflect the actual physiological activity situation. In the experiments, the SNR values of real and reconstructed BOLD signals were shown in Table 2. From the results, it could be seen that the signal-to-noise ratios in the two states were improved except for fmriNF, the motion imagery state. This was also consistent with the results demonstrated by t-SNE above, i.e., the results were worse in the fmriNF state and better in the other two states.
Figure 6: Visualization of real BOLD signal points with low time resolution and synthetic BOLD signal curves with high time resolution. The graph shows that the synthetic signal has the same trend as the real signal.
Data visualization was also performed. In the parietal region, where the correlation of motor imagery was strong, a node at the cortex corresponding to the C3 electrode portion of the EEG international standard lead system was selected. Curve graphs were used to display the signal intensities of the reconstructed high temporal-resolution time-series signals and the real signals at this node at the corresponding time points. As shown in Fig.6, the reconstructed high temporal resolution signal and the real signal exhibited a high degree of similarity in trend. This further suggests that the hyper-temporal resolution reconstruction results obtained using the CATD framework can effectively reflect the trend of blood oxygenation.
3.4 Evaluation of the effect of temporal features on the reconstructed signals
Different frequency bands of EEG data reflect various types of brain activity, and these differences can significantly impact the results. The effect of EEG frequency bands on BOLD signal reconstruction results was investigated. The EEG signals of different frequency bands were obtained by band-pass filtering, and BOLD signal reconstruction was performed using these signals. As shown in Fig.3(b), the beta and gamma bands, which are related to motor imagery, performed better in the classification index. Additionally, the timing metrics of the reconstruction results were calculated. As shown in Fig.7, the gamma band performed well in the timing metrics, while the beta band showed lower timing metrics.
Figure 7: Cosine similarity and CCC comparison of the timing of synthesized BOLD signals with real BOLD signals using EEG signals from different frequency bands.
The reasons for these results are analysed below. The brain’s state changes in preparation for or during imagined movements are generally slower and smoother, leading to event-related desynchronization (ERD) in the beta band[40]. This phenomenon may explain why the beta band excels in motor imagery and resting state categorization, even surpassing the full band’s categorization performance. However, this characteristic also causes the reconstructed results of the beta band to differ considerably in temporal similarity, resulting in a lower temporal index compared to the blood oxygenation activity reflected by the full-band brain electrical activity, i.e., the real BOLD signal. On the other hand, the gamma band is closely associated with higher cognitive and perceptual functions. In motor imagery tasks, the brain’s requirement to understand instructions and make judgments involves higher cognitive and perceptual functions[41]. Therefore, the gamma band most closely matches the real results in both the categorization index and timing index. The other three low-frequency bands: delta, theta, and alpha, performed poorly on both the classification metric (e.g., Fig.3(b)) and the temporal similarity metric (e.g., Fig.7), suggesting that these bands negatively affect the reconstruction results by containing less information about the correlation between the EEG and the BOLD signal.
3.5 Support for medical decision-making
Figure 8: Comparison of difference maps of BOLD function maps synthesized using EEG from two healthy subjects and Parkinson’s patients, with the portion marked by the red circle being the area of the brain that embodies the abnormality in both difference maps.
To demonstrate the potential application of the proposed CATD framework in the medical domain, cross-modal reconstruction experiments were conducted on Parkinson’s patients to support medical decision-making. In this part of the experiment, the framework was first trained on the NODDI dataset containing paired EEG and BOLD fMRI data from healthy subjects at resting state. Subsequently, another dataset containing only resting-state EEG data from Parkinson’s patients was used to reconstruct the cortical BOLD functional maps of these patients using the trained cross-modal generation model.
Predictions of Parkinson’s disease were performed using the true EEG signal, the reconstructed BOLD signal, and a combination of the true EEG signal and the reconstructed BOLD signal. The classification metrics are shown in Fig.3(c). Results indicate that the simultaneous use of reconstructed BOLD signals and recorded EEG signals significantly improves the accuracy of Parkinson’s disease prediction and related metrics. This suggests that the CATD framework effectively captures the potential connection between BOLD signals and EEG signals and can be applied across different datasets, which is crucial for disease diagnosis. Consequently, the framework not only enhances the accuracy of existing diagnostic methods but also provides new tools and methodologies for the diagnosis and research of other neurological diseases.
In order to fully utilize the advantage of our CATD framework, i.e., to obtain BOLD signals with high spatial resolution without the condition of fMRI detection, we performed a difference analysis of reconstructed BOLD functional maps in two healthy subjects and two Parkinson’s patients. As shown in Fig.8, both showed significant abnormalities in the brain regions marked by red circles. This region is the lingual gyrus, the medial occipito-temporal gyrus and the lateral occipito-temporal gyrus[42], which, in patients with Parkinson’s disease, shows significant structural changes[43], and in patients with PD accompanied by visual hallucinations, the atrophy of these three brain regions correlates with the severity of visual hallucinations[44]. This result demonstrates the potential application of the proposed method in localizing regions of abnormal brain activity and further proves its practical value in medical diagnosis.
4 Discussion
Cross-modal neuroimage synthesis is becoming crucial in neuroanalysis research. For the first time, we propose the CATD framework based on diffusion models to achieve cross-modal synthesis of temporal functional neuroimages, enabling the conversion of EEG to BOLD signals. This approach addresses limitations in BOLD acquisition, such as the inability of patients with metal implants to undergo fMRI scans. Enhancing BOLD temporal resolution facilitates the study of sub-millimeter cortical structures and activities[45], improves the quality and reliability of functional connectivity and task-driven fMRI research through higher sampling rates[46], and allows for more accurate detection and localization of functional activation regions, capturing transient neural activities[47]. This work provides novel insights for functional neuroimage synthesis[48]. Next, we will analyze the experimental results in detail to illustrate the performance of our framework.
In the experiments on the motor imagery dataset, our CATD framework demonstrates its superior performance in cross-modal feature learning and synthesis by detecting categorical metrics for both motor imagery states and resting states. Specifically, the framework is able to capture the features of EEG and BOLD signals well and perform effective cross-modal synthesis. This improved performance can be attributed to the ability to capture complex modes and integrate different modal features in the CATD framework, which enables it to achieve accurate transitions between various signal features. Ablation experiments further validate the effectiveness of the proposed CAB in aligning high-dimensional mismatched data pairs. CAB can learn and capture valid links between EEG and BOLD signals, ensuring the robustness of the model in cross-modal feature capture and synthesis.
In the temporal super-resolution experiments, the results exhibit a high correlation in all three motion imagery states, while the signal-to-noise ratio is improved in both states compared to the original signal. This suggests that temporal resolution enhancement is indeed feasible and can provide a more detailed and accurate temporal representation of the BOLD signal.
In order to further verify which EEG frequency bands have a greater impact on the results and have the potential to improve the efficiency of cross-modal synthesis, a frequency band analysis was performed. It was found that the beta band had the greatest impact on the classification results, while the gamma band performed the best in terms of similarity to the real signal. This can be explained by the properties of EEG signals. Lower frequency bands such as beta show event-related desynchronization (ERD) phenomenon in motor imagery tasks and therefore perform better in classification metrics. Whereas the higher cognitive functions involved in the motor imagery task are mainly determined by the gamma band, the reconstruction results of the gamma band therefore show the highest similarity to the original signal.
The potential of the proposed method to enhance the accuracy of disease diagnosis was validated through disease decision support experiments. The results demonstrate that the method is capable of identifying potentially abnormal regions in the brain by leveraging the high spatial resolution of synthetic BOLD signals, even when only EEG signals are available. This capability contributes to improved diagnostic accuracy.
While the results are promising, achieving accurate mapping from EEG to BOLD signals still presents challenges. Fig.9 illustrates a typical failure case, where the lower regions of the reconstructed BOLD map significantly differ from the real image. This discrepancy is likely due to the low spatial resolution and relatively low signal-to-noise ratio (SNR) of EEG signals. Even though our model successfully learns the relationship between EEG and BOLD signals, the process of upsampling to generate high-resolution images introduces inaccuracies. Addressing these issues of spatial resolution and SNR will be crucial for enhancing the precision of cross-modal synthesis in future work.
Figure 9: Typical failure case of our method. It can be seen that regions such as the lower part of the BOLD function map (near the frontal lobe) are more differentiated
Although the limited number of paired data subjects constrained the training, the amount of training data was increased by sampling and pairing the first 6 seconds of each frame of EEG data with a functional map of the BOLD signal through a data segmentation technique. This data augmentation approach positively impacted the model’s performance. All experiments were conducted across different subjects. While the results are not yet perfect, they still indicate a more favorable outcome.
5 Conclusion
In this work, a novel CATD framework is proposed for the cross-modal conversion of functional neuroimages, specifically the synthesis of BOLD signals from EEG signals. To fully exploit the high temporal resolution of EEG signals, the DTFS module was designed to increase the sampling rate of the EEG signal as a conditioned signal, achieving temporal resolution enhancement of the synthesized BOLD signal. By constructing the CAB module, the alignment of high-dimensional heterogeneous functional neuroimages in the hidden space was realized. Qualitative and quantitative experimental results demonstrate that the proposed framework effectively achieves cross-modal synthesis from EEG to BOLD signals. The effectiveness of CAB was validated through ablation experiments, and the framework’s value was illustrated in practical application scenarios through medical decision support experiments. Future studies will focus on further optimizing the model and improving the quality of the reconstructed signals to achieve more comprehensive functional neuroimage synthesis.
Acknowledgment